### v1

# Jumping Automata Must Pay

Abstract
Jumping automata are finite automata that read their input in a non-sequential manner, by allowing a reading head to “jump” between positions on the input, consuming a permutation of the input word. We argue that allowing the head to jump should incur some cost. To this end, we propose three quantitative semantics for jumping automata, whereby the jumps of the head in an accepting run define the cost of the run. The three semantics correspond to different interpretations of jumps: the absolute distance semantics counts the distance the head jumps, the reversal semantics counts the number of times the head changes direction, and the Hamming distance measures the number of letter-swaps the run makes.
We study these measures, with the main focus being the boundedness problem: given a jumping automaton, decide whether its (quantitative) languages is bounded by some given number $k$. We establish the decidability and complexity for this problem under several variants.
1 Introduction
Traditional automata read their input sequentially. Indeed, this is the case for most state-based computational models. In some settings, however, we wish to abstract away the order of the input letters. For example, when the input represents available resources, and we only wish to reason about their quantity. From a more language-theoretic perspective, this amounts to looking at the commutative closure of languages, a.k.a. their Parikh image. To capture this notion in a computation model, Jumping Automata (JFAs) were introduced in [20]. A jumping automaton may read its input in a non-sequential manner, jumping from letter to letter, as long as every letter is read exactly once. Several works have studied the algorithmic properties and expressive power of these automata [12, 13, 22, 11, 18, 5].
While JFAs are an attractive and simple model, they present a shortcoming when thought of as model for systems, namely that the abstraction of the order may be too coarse. More precisely, the movement of the head can be thought of as a physical process of accessing the input storage of the JFA. Then, sequential access is the most basic form of access and can be considered “cheap”, but allowing the head to jump around is physically more difficult and therefore should incur a cost.
To address this, we present three quantitative semantics for JFAs, whereby a JFA represents a function from words to costs, which captures how expensive it is to accept a given word with respect to the head jumps. The different semantics capture different aspects of the cost of jumps, as follows.
Consider a JFA $\mathcal{A}$ and a word $w$, and let $\rho$ be an accepting run of $\mathcal{A}$ on $w$. The run $\rho$ specifies the sequence of states and indices visited in $w$. We first define the cost of individual runs.
  • In the Absolute Distance semantics (abs), the cost of $\rho$ is the sum of the lengths of jumps it makes.
  • In the Reversal semantics (rev), the cost of $\rho$ is the number of times the reading head changes its direction (i.e., moving from right to left or from left to right).
  • In the Hamming semantics (ham), we consider the word $w^{\prime}$ induced by $\rho$, i.e., the order of letters that $\rho$ reads. Then, the cost of $\rho$ is the number of letters where $w^{\prime}$ differs from $w$.
We then define the cost of the word $w$ in $\mathcal{A}$ according to each semantics, by taking the run that minimizes the cost.
Thus, we lift JFAs from a Boolean automata model to the rich setting of quantitative models [6, 4, 9, 10]. Unlike other quantitative automata, however, the semantics in this setting arise naturally from the model, without an external domain. Moreover, the definitions are naturally motivated by different types of memory access, as we now demonstrate. First, consider a system whose memory is laid out in an array (i.e., a tape), with a reading head that can move along the tape. Moving the head requires some energy, and therefore the total energy spent reading the input corresponds to the abs semantics. Next, consider a system whose memory is a spinning disk (or a sliding tape), so that the head stays in place and the movement is of the memory medium. Then, it is cheap to continue spinning in the same direction [^1] We assume that the head does not return back to the start by continuing the spin, but rather reaches some end., and the main cost is in reversing the direction, which requires stopping and reversing a motor. Then, the rev semantics best captures the cost. Finally, consider a system that reads its input sequentially, but is allowed to edit its input by replacing one letter with another, such that at the end the obtained word is a permutation of the original word. This is akin to edit-distance automata [21, 14] under a restriction of maintaining the amount of resources. Then, the minimal edits required correspond to the ham semantics.
Example 1.
Consider a (standard) NFA $\mathcal{A}$ for the language given by the regular expression $(ab)^{*}$ (the concrete automaton chosen is irrelevant, see Remark 6). As a JFA, $\mathcal{A}$ accepts a word $w$ if and only if $w$ has an equal number of $a$’s and $b$’s. To illustrate the different semantics, consider the words $w_{1}=ababbaab$ and $w_{2}=ababbaba$, obtained from $(ab)^{4}$ by flipping the third $ab$ pair ($w_{1}$) and the third and fourth pairs ($w_{2}$). As we define in Section 3, we think of runs of the JFA $\mathcal{A}$ as if the input is given between end markers at indices 00 and $n+1$, and the jumping run must start at 00 and end in $n+1$.
  • In the abs semantics, the cost of $w_{1}$, denoted $\mathcal{A}_{\textsc{abs}}(w_{1})$, is $2$: the indices read by the head are $0,1,2,3,4,6,5,7,8,9$, so there are two jumps of cost $1$: from $4$ to $6$ and from $5$ to $7$. Similarly, we have $\mathcal{A}_{\textsc{abs}}(w_{2})=4$, e.g., by the sequence $0,1,2,3,4,6,5,8,7,9$, which has two jumps of cost $1$ ($4$ to $6$ and $7$ to $9$), and one jump of cost $2$ ($5$ to $8$). (formally, we need to prove that there is no better run, but this is not hard to see).
  • In the rev semantics, we have $\mathcal{A}_{\textsc{rev}}(w_{1})=2$ by the same sequence of indices as above, as the head performs two “turns”, one at index $6$ (from $\to$ to $\leftarrow$) and one at $5$ (from $\leftarrow$ to $\to$). Here, however, we also have $\mathcal{A}_{\textsc{rev}}(w_{2})=2$, using the sequence $0,1,2,3,4,6,8,7,5,9$, whose turning points are $8$ and $5$.
  • In the ham semantics we have $\mathcal{A}_{\textsc{ham}}(w_{1})=2$ and $\mathcal{A}_{\textsc{ham}}(w_{2})=4$, since we must change the letters in all the flipped pairs for the words to be accepted.
Example 2.
Consider now an NFA $\mathcal{B}$ for the language given by the regular expression $a^{*}b^{*}$. Note that as a JFA, $\mathcal{B}$ accepts $\{a,b\}^{*}$, since every word can be reordered to the form $a^{*}b^{*}$.
Observe that in the rev semantics, for every word $w$ we have $\mathcal{B}_{\textsc{rev}}(w)\leq 2$, since at the worst case $\mathcal{B}$ makes one left-to-right pass to read all the $a$’s, then a right-to-left pass to read all the $b$’s, and then jump to the right end marker, and thus it has two turning points. In particular, $\mathcal{B}_{\textsc{rev}}$ is bounded.
However, in the abs and ham semantics, the costs can become unbounded. Indeed, in order to accept words of the form $b^{n}a^{n}$, in the abs semantics the head must first jump over all the $b^{n}$, incurring a high cost, and for the ham semantics, all the letters must be changed, again incurring a high cost.
Related work
Jumping automata were introduced in [20]. We remark that [20] contains some erroneous proofs (e.g., closure under intersection and complement, also pointed out in [13]). The works in [12, 13] establish several expressiveness results on jumping automata, as well as some complexity results. In [22] many additional closure properties are established. An extension of jumping automata with a two-way tape was studied in [11], and jumping automata over infinite words were studied by the first author in [5].
When viewed as the commutative image of a language, jumping automata are closely related to Parikh Automata [17, 8, 7, 15], which read their input and accept if a certain Parikh image relating to the run belongs to a given semilinear set (indeed, we utilize the latter in our proofs). Another related model is that of symmetric transducers – automata equipped with outputs, such that permutations in the input correspond to permutations in the output. These were studied in [3] in a jumping-flavour, and in [2] in a quantitative $k$-window flavour.
More broadly, quantitative semantics have received much attention in the past two decades, with many motivations and different flavors of technicalities. We refer the reader to [6, 10] and the references therein.
Contribution and paper organization
Our contribution consists of the introduction of the three jumping semantics, and the study of decision problems pertaining to them (defined in Section 3). Our main focus is the boundedness problem: given a JFA $\mathcal{A}$, decide whether the function described by it under each of the semantics bounded by some constant $k$. We establish the decidability of this problem for all the semantics, and consider the complexity of some fragments.
Our paper is organized as follows: the preliminaries and definitions are given in Sections 2 and 3. Then, each of Sections 4, 5 and 6 studies one of the semantics, and follows the same structure: we initially establish that the membership problem for the semantics is NP-complete. Then we characterize the set of words whose cost is at most $k$ using a construction of an NFA. These constructions differ according to the semantics, and involve some nice tricks with automata, but are technically not hard to understand. We note that these constructions are preceded by crucial observations regarding the semantics, which allow us to establish their correctness. Next, in Section 7 we give a complete picture of the interplay between the different semantics (using some of the results established beforehand). Finally, in Section 8 we discuss some exciting open problems. Due to lack of space, some proofs appear in the appendix.
2 Preliminaries and Definitions
For a finite alphabet $\Sigma$ we denote by $\Sigma^{*}$ the set of finite words over $\Sigma$. For $w\in\Sigma^{*}$ we denote its letters by $w=w_{1}\cdots w_{n}$, and its length by $|w|=n$. In the following, when discussing sets of numbers, we define $\min\emptyset=\infty$
Automata
A nondeterministic finite automaton (NFA) is a 5-tuple $\mathcal{A}=\langle\Sigma,Q,\delta,Q_{0},\alpha\rangle$ where $\Sigma$ is a finite alphabet, $Q$ is a finite set of states, $\delta:Q\times\Sigma\to 2^{Q}$ is a nondeterministic transition function, $Q_{0}\subseteq Q$ is a set of initial states, and $\alpha\subseteq Q$ is a set of accepting states. A run of $\mathcal{A}$ on a word $w=w_{1}w_{2}\ldots w_{n}$ is a sequence $\rho=q_{0},q_{1},\ldots,q_{n}$ such that $q_{0}\in Q_{0}$ and for every $0\leq i<n$ it holds that $q_{i+1}\in\delta(q_{i},w_{i+1})$. The run $\rho$ is accepting if $q_{n}\in\alpha$. A word $w$ is accepted by $\mathcal{A}$ if there exists an accepting run of $\mathcal{A}$ on $w$. The language of $\mathcal{A}$, denoted $\mathfrak{L}(\mathcal{A})$, is the set of words accepted by $\mathcal{A}$.
Permutations
Let $n\in\mathbb{N}$. The permutation group $S_{n}$ is the set of bijections (i.e. permutations) from $\{1,...,n\}$ to itself. $S_{n}$ forms a group with the function-composition operation and the identity permutation as a neutral element. Given a word $w=w_{1}\cdots w_{n}$ and a permutation $\pi\in S_{n}$, we define $\pi(w)=w_{\pi(1)}\cdots w_{\pi(n)}$. For example, if $w=abcd$ and $\pi=\begin{pmatrix}1&2&3&4\\ 3&4&2&1\end{pmatrix}$ then $\pi(w)=cdba$. We usually denote permutations in one-line form, e.g., $\pi$ is represented as $(3,4,2,1)$. We say that a word $y$ is a permutation of $x$, and we write $x\sim y$ if there exists a permutation $\pi\in S_{|x|}$ such that $\pi(x)=y$.
Jumping Automata
A jumping automaton is syntactically identical to an NFA, with the semantic difference that it has a reading head that can “jump” between indices of the input word. An equivalent view is that a jumping automaton reads a (nondeterministically chosen) permutation of the input word.
Formally, consider an NFA $\mathcal{A}$. We view $\mathcal{A}$ as a jumping finite automaton (JFA) by defining its jumping language $\mathfrak{J}(\mathcal{A})=\{w\in\Sigma^{*}\mid\exists u\in\Sigma^{*}.\ w\sim u\wedge u\in\mathfrak{L}(\mathcal{A})\}$.
Since our aim is to reason about the manner with which the head of a JFA jumps, we introduce a notion to track the head along a run. Consider a word $w$ of length $n$ and a JFA $\mathcal{A}$. A jump sequence is a vector $\mathbf{a}=(a_{0},a_{1},a_{2},\ldots,a_{n},a_{n+1})$ where $a_{0}=0$, $a_{n+1}=n+1$ and $(a_{1},a_{2},\ldots,a_{n})\in S_{n}$. We denote by $J_{n}$ the set of all jump sequences of size $n+2$.
Intuitively, a jump sequence $\mathbf{a}$ represents the order in which a JFA visits a given word of length $n$. First it visits the letter at index $a_{1}$, then the letter at index $a_{2}$ and so on. To capture this, we define $w_{\mathbf{a}}=w_{a_{1}}w_{a_{2}}\cdots w_{a_{n}}$. Observe that jump sequences enforce that the head starts at position 0 and ends at position $n+1$, which can be thought of as left and right markers, as is common in e.g., two-way automata.
An alternative view of jumping automata is via Parikh Automata (PA) [17, 7]. The standard definition of PA is an automaton whose acceptance condition includes a semilinear set over the transitions. To simplify things, and to avoid defining unnecessary concepts (e.g., semilinear sets), for our purposes, a PA is a pair $(\mathcal{A},\mathcal{C})$ where $\mathcal{A}$ is an NFA over alphabet $\Sigma$, and $\mathcal{C}$ is a JFA over $\Sigma$. Then, the PA $(\mathcal{A},\mathcal{C})$ accepts a word $w$ if $w\in\mathfrak{L}(\mathcal{A})\cap\mathfrak{J}(\mathcal{C})$. Note that when $\mathfrak{L}(\mathcal{A})=\Sigma^{*}$, then the PA coincides with $\mathfrak{J}(\mathcal{C})$. Our usage of PA is to obtain the decidability of certain problems. Specifically, from [17] we have that emptiness of PA is decidable.
3 Quantitative Semantics for JFAs
In this section we present and demonstrate the three quantitative semantics for JFAs. We then define the relevant decision problems, and lay down some general outlines to solving them, which are used in later sections. For the remainder of the section fix a JFA $\mathcal{A}=\langle\Sigma,Q,\delta,Q_{0},\alpha\rangle$.
3.1 The Semantics
The Absolute-Distance Semantics
In the absolute-distance semantics, the cost of a run (given as a jump sequence) is the sum of the sizes of the jumps made by the head. Since we want to think of a sequential run as a run with 0 jumps, we measure a jump over $k$ letters as distance $k-1$ (either to the left or to the right). This is captured as follows.
For $k\in\mathbb{Z}$, define $\llbracket k\rrbracket=|k|-1$. Consider a word $w\in\Sigma^{*}$ with $|w|=n$, and let $\mathbf{a}=(a_{0},a_{1},a_{2},\ldots,a_{n},a_{n+1})$ be a jump sequence, then we lift the notation above and write $\llbracket\mathbf{a}\rrbracket=\sum_{i=1}^{n+1}\llbracket a_{i}-a_{i-1}\rrbracket$.
Definition 3 (Absolute-Distance Semantics).
For a word $w\in\Sigma^{*}$ with $|w|=n$ we define
$\mathcal{A}_{\textsc{abs}}(w)=\min\{\llbracket\mathbf{a}\rrbracket\mid\mathbf{a}\text{ is a jump sequence, and }w_{\mathbf{a}}\in\mathfrak{L}(\mathcal{A})\}$
(recall that $\min\emptyset=\infty$ by definition).
The Reversal Semantics
In the reversal semantics, the cost of a run is the number of times the head changes direction in the corresponding jump sequence. Consider a word $w\in\Sigma^{*}$ with $|w|=n$, and let $\mathbf{a}=(a_{0},a_{1},a_{2},\ldots,a_{n},a_{n+1})$ be a jump sequence, we define
#$\#_{\textsc{rev}}(\mathbf{a})=|\{i\in\{1,\ldots,n\}\mid(a_{i}>a_{i-1}\wedge a_{i}>a_{i+1})\vee(a_{i}<a_{i-1}\wedge a_{i}<a_{i+1})\}|$
Definition 4 (Reversal Semantics).
For a word $w\in\Sigma^{*}$ with $|w|=n$ we define
$\mathcal{A}_{\textsc{rev}}(w)=\min\{\#_{\textsc{rev}}(\mathbf{a})\mid\mathbf{a}\text{ is a jump sequence, and }w_{\mathbf{a}}\in\mathfrak{L}(\mathcal{A})\}$
The Hamming Semantics
In the Hamming measure, the cost of a word is the minimal number of coordinates of $w$ that need to be changed in order for the obtained word to be accepted by $\mathcal{A}$ (sequentially, as an NFA), so that the changed word is a permutation of $w$.
Consider two words $x,y\in\Sigma^{*}$ with $|x|=|y|=n$ such that $x\sim y$, we define the Hamming Distance between $x$ and $y$ as $d_{H}(x,y)=|\{i\mid x_{i}\neq y_{i}\}|$.
Definition 5 (Hamming Semantics).
For a word $w\in\Sigma^{*}$ we define
$\mathcal{A}_{\textsc{ham}}(w)=\min\{d_{H}(w^{\prime},w)\mid w^{\prime}\in L(\mathcal{A}),w^{\prime}\sim w\}$
Remark 6.
Note that the definitions of the three semantics are independent of the NFA, and only refer to its language. We can therefore refer to the cost of a word in a language according to each semantics, rather than the cost of a word in a concrete automaton.
3.2 Quantitative Decision Problems
In the remainder of the paper we focus on quantitative variants of the standard Boolean decision problems pertaining to the jumping semantics. Specifically, we consider the following problems for each semantics $\textsc{sem}\in\{\textsc{abs},\textsc{ham},\textsc{rev}\}$.
  • Membership: Given a JFA $\mathcal{A}$, $k\in\mathbb{N}$ and a word $w$, decide whether $\mathcal{A}_{\textsc{sem}}(w)\leq k$.
  • $k$-Boundedness (for a fixed $k$): Given a JFA $\mathcal{A}$, decide whether $\forall w\in\mathfrak{J}(\mathcal{A})\ \mathcal{A}_{\textsc{sem}}(w)\leq k$.
  • Param-Boundedness: Given a JFA $\mathcal{A}$ and $k\in\mathbb{N}$, decide whether $\forall w\in\mathfrak{J}(\mathcal{A})\ \mathcal{A}_{\textsc{sem}}(w)\leq k$.
We also pay special attention to the setting where $\mathfrak{J}(\mathcal{A})=\Sigma^{*}$, in which case we refer to these problems as Univ-$k$-Boundedness and Univ-Param-Boundedness. For example, in Univ-Param-Boundedness we are given a JFA $\mathcal{A}$ and $k\in\mathbb{N}$ and the problem is to decide whether $\mathcal{A}_{\textsc{sem}}(w)\leq k$ for all words $w\in\Sigma^{*}$.
The boundedness problems can be thought of as quantitative variants of Boolean universality (i.e., is the language equal to $\Sigma^{*}$). Observe that the problems above are not fully specified, as the encoding of $k$ (binary or unary) when it is part of the input may effect the complexity. We remark on this when it is relevant. Note that the emptiness problem is absent from the list above. Indeed, a natural quantitative variant would be: is there a word $w$ such that $\mathcal{A}_{\textsc{sem}}(w)\leq k$. This, however, is identical to Boolean emptiness, since $\mathfrak{L}(\mathcal{A})\neq\emptyset$ if and only if there exists $w$ such that $\mathcal{A}_{\textsc{sem}}(w)=0$. We therefore do not consider this problem. Another problem to consider is boundedness when $k$ is existentially quantified. We elaborate on this problem in Section 8.
4 The Absolute-Distance Semantics
The first semantics we investigate is abs, and we start by showing that (the decision version of) computing its value for a given word is NP-complete. This is based on bounding the distance with which a word can be accepted.
Lemma 7.
Consider a JFA $\mathcal{A}$ and $w\in\mathfrak{J}(\mathcal{A})$ with $|w|=n$, then $\mathcal{A}_{\textsc{abs}}(w)<n^{2}$.
Proof.
Since $w\in\mathfrak{J}(\mathcal{A})$, there exists a jump sequence $\mathbf{a}=(a_{0},a_{1},a_{2},\ldots,a_{n},a_{n+1})$ such that $w_{\mathbf{a}}\in\mathfrak{L}(\mathcal{A})$. Therefore, $\mathcal{A}_{\textsc{abs}}(w)\leq\llbracket\mathbf{a}\rrbracket$. Observe that $|a_{i}-a_{i-1}|\leq{n}$ for all $i\in\{1,\ldots,n+1\}$, since there is no jump from 00 to $n+1$ (since $a_{0}=0$ and $a_{n}=n+1$). The following concludes the proof:
$\llbracket\mathbf{a}\rrbracket=\sum_{i=1}^{n+1}\llbracket a_{i}-a_{i-1}\rrbracket=\sum_{i=1}^{n+1}|a_{i}-a_{i-1}|-1\leq\sum_{i=1}^{n+1}n-1=(n+1)(n-1)<n^{2}$
We can now prove the complexity bound for computing the absolute distance, as follows.
Theorem 8 (Absolute-Distance Membership is NP-complete).
The problem of deciding, given $\mathcal{A},w$ and $k\in\mathbb{N}$, whether $\mathcal{A}_{\textsc{abs}}(w)\leq k$, is NP-complete.
Proof.
In order to establish membership in NP, note that by Lemma 7, we can assume $k\leq n^{2}$, as otherwise we can set $k=n^{2}$. Then, it is sufficient to nondeterministically guess a jump sequence $\mathbf{a}$ and to check that $w_{a_{1}}\cdots w_{a_{n}}\in\mathfrak{L}(\mathcal{A})$ and that $\llbracket\mathbf{a}\rrbracket\leq k$. Both conditions are easily checked in polynomial time, since $k$ is polynomially bounded.
Hardness in NP follows by reduction from (Boolean) membership in JFA: it is shown in [13] that deciding whether $w\in\mathfrak{J}(\mathcal{A})$ is NP-hard. We reduce this problem by outputting, given $\mathcal{A}$ and $w$, the same $\mathcal{A}$ and $w$ with the bound $k=n^{2}$. The reduction is correct by Lemma 7 and the fact that if $w\notin\mathfrak{J}(\mathcal{A})$ then $\mathcal{A}_{\textsc{abs}}(w)=\infty$. ∎
4.1 Decidability of Boundedness Problems for abs
We now turn our attention to the boundedness problems. Consider a JFA $\mathcal{A}$ and $k\in\mathbb{N}$. Intuitively, our approach is to construct an NFA $\mathcal{B}$ that simulates, while reading a word $w\in\Sigma^{*}$, every jump sequence of $\mathcal{A}$ on $w$ whose absolute distance is at most $k$. The crux of the proof is to show that we can indeed bound the size of $\mathcal{B}$ as a function of $k$. At a glance, the main idea here is to claim that since the absolute distance is bounded by $k$, then $\mathcal{A}$ cannot make large jumps, nor many small jumps. Then, if we track a sequential head going from left to right, then the jumping head must always be within a bounded distance from it. We now turn to the formal arguments. Fix a JFA $\mathcal{A}=\langle\Sigma,Q,\delta,Q_{0},\alpha\rangle$.
To understand the next lemma, imagine $\mathcal{A}$’s jumping head while taking the $j^{th}$ step in a run on $w$ according to a jump sequence $\mathbf{a}=(a_{0},a_{1},a_{2},\ldots,a_{n},a_{n+1})$. Thus, the jumping head points to the letter at index $a_{j}$. Concurrently, imagine a “sequential” head (reading from left to right), which points to the $j^{th}$ letter in $w$. Note that these two heads start and finish reading the word at the same indices $a_{0}=0$ and $a_{n+1}=n+1$. It stands to reason that if at any step while reading $w$ the distance between these two heads is large, the cost of reading $w$ according to $\mathbf{a}$ would also be large, as there would need to be jumps that bridge the gaps between the heads. The following lemma formalizes this idea.
Lemma 9.
Consider a jump sequence $\mathbf{a}=(a_{0},a_{1},a_{2},\ldots,a_{n},a_{n+1})$. For every $1\leq j\leq n$ it holds that $\llbracket\mathbf{a}\rrbracket\geq|a_{j}-j|$.
Proof.
[…]
$\left(\sum_{i=1}^{j}|a_{i}-a_{i-1}|\right)-j\underset{\mathclap{\begin{subarray}{c}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\big{\downarrow}\\ {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{triangle inequality}}\end{subarray}}}{\geq}\left|\sum_{i=1}^{j}a_{i}-a_{i-1}\right|-j\underset{\mathclap{\begin{subarray}{c}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\big{\downarrow}\\ {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{telescopic sum}}\end{subarray}}}{=}|a_{j}-a_{0}|-j\underset{\mathclap{\begin{subarray}{c}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\big{\downarrow}\\ {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{0}=0}\end{subarray}}}{=}a_{j}-j\underset{\mathclap{\begin{subarray}{c}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\big{\downarrow}\\ {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{j}\geq j}\end{subarray}}}{=}|a_{j}-j|$
The direction $a_{j}<j$ is proved by looking at the sum of the last $j$ elements (see in Section A.1). ∎
From Lemma 9 we get that in order for a word $w$ to attain a small cost, it must be accepted with a jumping sequence that stays close to the sequential head. More precisely:
Corollary 10.
Let $k\in\mathbb{N}$ and consider a word $w$ such that $\mathcal{A}_{\textsc{abs}}(w)\leq k$, then there exists a jumping sequence $\mathbf{a}=(a_{0},a_{1},a_{2},\ldots,a_{n},a_{n+1})$ such that $w_{\mathbf{a}}\in\mathfrak{L}(\mathcal{A})$ and for all $1\leq j\leq n$ we have $|a_{j}-j|\leq k$.
We now turn to the construction of an NFA that recognizes the words whose cost is at most $k$.
Lemma 11.
Let $k\in\mathbb{N}$. We can effectively construct an NFA $\mathcal{B}$ such that $\mathfrak{L}(\mathcal{B})\!=\!\{w\in\Sigma^{*}|\mathcal{A}_{\textsc{abs}}(w)\leq k\}$.
Proof.
Let $k\in\mathbb{N}$. Intuitively, $\mathcal{B}$ works as follows: it remembers in its states a window of size $12k+$ 1 centered around the current letter (recall that as an NFA, it reads its input sequentially). The window is constructed by nondeterministically guessing (and then verifying) the next $k$ letters, and remembering the last $k$ letters.
$\mathcal{B}$ then nondeterministically simulates a jumping sequence of $\mathcal{A}$ on the given word, with the property that the jumping head stays within distance $k$ from the sequential head. This is done by marking for each letter in the window whether it has already been read in the jumping sequence, and nondeterministically guessing the next letter to read, while keeping track of the current jumping head location, as well as the total cost incurred so far. After reading a letter, the window is shifted by one to the right. If at any point the window is shifted so that a letter that has not been read by the jumping head shifts out of the $12k+$ 1 scope, the run rejects. Similarly, if the word ends but the guessed run tried to read a letter beyond the length of the word, the run rejects. The correctness of the construction follows from Corollary 10. We now turn to the formal details. Recall that $\mathcal{A}=\langle\Sigma,Q,\delta,Q_{0},\alpha\rangle$. We define $\mathcal{B}=\langle\Sigma,Q^{\prime},\delta^{\prime},Q_{0}^{\prime},\beta\rangle$ as follows.
The state space of $\mathcal{B}$ is $Q^{\prime}=Q\times(\Sigma\times\{?,\checkmark\})^{-k,\ldots,k}\times\{-k,\ldots,k\}\times\{0,\ldots,k\}$. We denote a state of $\mathcal{B}$ as $(q,f,j,c)$ where $q\in Q$ is a state of $\mathcal{A}$, $f:\{-k,\ldots,k\}\to\Sigma\times\{?,\checkmark\}$ represents a window of size $12k+$ 1 around the sequential head, where $\checkmark$ marks letters that have already been read by $\mathcal{A}$ (and ???? marks the others), $j$ represents the index of the head of $\mathcal{A}$ relative to the sequential head, and $c$ represents the cost incurred thus far in the run. We refer to the components of $f$ as $f(j)=(f(j)_{1},f(j)_{2})$ with $f(j)_{1}\in\Sigma$ and $f(j)_{2}\in\{?,\checkmark\}$.
The initial states of $\mathcal{B}$ are $Q^{\prime}_{0}=\left\{(q,f,j,j-1)\mid q\in Q_{0}\wedge j>0\wedge(f(i)_{2}=\checkmark\iff i\leq 0)\right\}$. That is, all states where the state of $\mathcal{A}$ is initial, the location of the jumping head is some $j>0$ incurring a cost of $j-1$ (i.e., the initial jump $\mathcal{A}$ makes), and the window is guessed so that everything left of the first letter is marked as already-read (to simulate the fact that $\mathcal{A}$ cannot jump to the left of the first letter).
The transitions of $\mathcal{B}$ are defined as follows. Consider a state $(q,f,j,c)$ and a letter $\sigma\in\Sigma$, then $(q^{\prime},f^{\prime},j^{\prime},c^{\prime})\in\delta^{\prime}((q,f,j,c),\sigma)$ if and only if the following hold (see Fig. 1 for an illustration):
  • $f(1)_{1}=\sigma$. That is, we verify that the next letter in the guessed window is indeed correct.
  • $f(-k)_{2}=\checkmark$. That is, the leftmost letter has been read. Otherwise by Corollary 10 the cost of continuing the run must be greater than $k$.
  • $f(j)_{2}\neq\checkmark$ and $f^{\prime}(j-1)=\checkmark$ (if $j>-k$). That is, the current letter has not been previously read, and will be read from now on (note that index $j$ before the transition corresponds to index $j-1$ after).
  • $q^{\prime}=\delta(q,f(j)_{1})$, i.e. the state of $\mathcal{A}$ is updated according to the current letter.
  • $c^{\prime}=c+|j^{\prime}+1-j|-1$, since $j^{\prime}$ represents the index in the shifted window, so in the “pre-shifted” tape this is actually index $j+1$. We demonstrate this in Fig. 1. Also, $c^{\prime}\leq k$ by the definition of $Q$.
  • $f^{\prime}(i)=f(i+1)$ for $i<k$. That is, the window is shifted and the index $f^{\prime}(k)$ is nondeterministically guessed [^2] The guess could potentially be $\checkmark$, but this is clearly useless..
$\alpha\checkmark$$\beta?$$\gamma\checkmark$$\epsilon?$$\mu\checkmark$$\eta?$$\xi?$$\zeta\checkmark$$\theta?$44-4- 433-3- 322-2- 211-1- 1001111222233334444$\implies$
$\beta\checkmark$$\gamma\checkmark$$\epsilon?$$\mu\checkmark$$\eta?$$\xi?$$\zeta\checkmark$$\theta?$$\psi?$44-4- 433-3- 322-2- 211-1- 1001111222233334444
Figure 1: A single transition in the construction of Lemma 11. The dashed arrow signifies the sequential head, the full arrow is the “imaginary” jumping head. Here, the head jumps from $-3$ to $2$, incurring a cost of $4$, but in the indexing after the transition $\xi$ is at index $1$, thus the expression given for $c^{\prime}$ in the construction. Note that the letter being read must be $\mu$, and that $\alpha$ must be checked, otherwise the run has failed.
Finally, the accepting states of $\mathcal{B}$ are $\beta=\{(q,f,1,c)\mid q\in\alpha\wedge f(j)_{2}=?\text{ for all }j>0\}$. That is, the state of $\mathcal{A}$ is accepting, the overall cost is at most $k$, the location of the jumping head matches the sequential head (intuitively, location $n+1$), and no letter beyond the end of the tape has been used.
It is easy to verify that $\mathcal{B}$ indeed guesses a jump sequence and a corresponding run of $\mathcal{A}$ on the given word, provided that the jumping head stays within distance $k$ of the sequential head. By Corollary 10, this restriction is complete, in the sense that if $\mathcal{A}_{\textsc{abs}}(w)\leq k$ then there is a suitable jump sequence under this restriction with which $w$ is accepted. ∎
We can now readily conclude the decidability of the boundedness problems for the abs semantics. The proof (in Section A.2) makes use of the decidability of emptiness for Parikh Automata [17].
Theorem 12.
The following problems are decidable for the abs semantics: $k$-Boundedness, Param-Boundedness, Univ-$k$-Boundedness and Univ-Param-Boundedness.
With further scrutiny, we see that the size of $\mathcal{B}$ constructed as per Lemma 11 is polynomial in the size of $\mathcal{A}$ and single-exponential in $k$. Thus, Univ-$k$-Boundedness is in fact decidable in PSPACE, whereas Univ-Param-Boundedness is in EXPSPACE and 2-EXPSPACE for $k$ given in unary and binary, respectively. For the non-universal problems we do not supply upper complexity bounds, as these depend on the decidability for PA containment, for which we only derive decidability from [17].
4.2 PSPACE-Hardness of Boundedness for abs
In the following, we complement the decidability result of Theorem 12 by showing that already Univ-$k$-Boundedness is PSPACE-hard, for every $k\in\mathbb{N}$.
We first observe that the absolute distance of every word is even. In fact, this is true for every jumping sequence.
Lemma 13.
Consider a jumping sequence $\mathbf{a}=(a_{0},a_{1},\ldots,a_{n},a_{n+1})$, then $\llbracket\mathbf{a}\rrbracket$ is even.
Proof.
Observe that the parity of $|a_{i}-a_{i-1}|$ is the same as that of $a_{i}-a_{i-1}$. It follows that the parity of $\llbracket\mathbf{a}\rrbracket=\sum_{i=1}^{n+1}\llbracket a_{i}-a_{i-1}\rrbracket=\sum_{i=1}^{n+1}|a_{i}-a_{i-1}|-1$ is the same as that of
$\sum_{i=1}^{n+1}(a_{i}-a_{i-1}-1)=\left(\sum_{i=1}^{n+1}a_{i}-a_{i-1}\right)-(n+1)=n+1-(n+1)=0$
and is therefore even (the penultimate equality is due to the telescopic sum). ∎
We say that $\mathcal{A}_{\textsc{abs}}$ is $k$-bounded if $\mathcal{A}_{\textsc{abs}}(w)\leq k$ for all $w\in\Sigma^{*}$. We are now ready to prove the hardness of Univ-$k$-Boundedness. Observe that for a word $w\in\Sigma^{*}$ we have that $\mathcal{A}_{\textsc{abs}}(w)=0$ if and only if $w\in\mathfrak{L}(\mathcal{A})$ (indeed, a cost of 00 implies that an accepting jump sequence is the sequential run $10,1,\ldots,|w|+$ 1). In particular, we have that $\mathcal{A}_{\textsc{abs}}$ is 00-bounded if and only if $\mathfrak{L}(\mathcal{A})=\Sigma^{*}$. Since the universality problem for NFAs is PSPACE-complete, this readily proves that Univ-00-Boundedness is PSPACE-hard. Note, however, that this does not imply that Univ-$k$-Boundedness is also PSPACE-hard for other values of $k$, and that the same argument fails for $k>0$. We therefore need a slightly more elaborate reduction.
Lemma 14.
For abs the Univ-$k$-Boundedness and $k$-Boundedness problems are PSPACE-hard for every $k\in\mathbb{N}$.
Proof.
We sketch the proof for Univ-$k$-Boundedness. The case of $k$-Boundedness requires slightly more effort and is delegated to Section A.3. By Lemma 13, we can assume without loss of generality that $k$ is even. Indeed, if there exists $m\in\mathbb{N}$ such that $\mathcal{A}_{\textsc{abs}}(w)\leq 2m+1$ for every $w\in\Sigma^{*}$, then by Lemma 13 we also have $\mathcal{A}_{\textsc{abs}}(w)\leq 2m$. Therefore, we assume $k=2m$ for some $m\in\mathbb{N}$.
We reduce the universality problem for NFAs to the Univ-$2m$-Boundedness problem. Consider an NFA $\mathcal{A}=\langle Q,\Sigma,\delta,Q_{0},\alpha\rangle$, and let $\heartsuit\notin\Sigma$ be a fresh symbol. Intuitively, we obtain from $\mathcal{A}$ an NFA $\mathcal{B}$ over the alphabet $\Sigma\cup\{\heartsuit\}$ such that $w\in\mathfrak{L}(\mathcal{B})$ if and only if the following hold:
  1. 1.
    Either $w$ does not contain exactly $m$ occurrences of $\heartsuit$, or
  2. 2.
    $w$ contains exactly $m$ occurrences of $\heartsuit$, but does not start with $\heartsuit$, and $w|_{\Sigma}\in\mathfrak{L}(\mathcal{A})$ (where $w|_{\Sigma}$ is obtained from $w$ by removing all occurrences of $\heartsuit$).
We then have the following: if $\mathfrak{L}(\mathcal{A})=\Sigma^{*}$, then for every $w\in(\Sigma\cup\{\heartsuit\})^{*}$ if $w\in\mathfrak{L}(\mathcal{B})$ then $\mathcal{B}_{\textsc{abs}}(w)=0\leq 2m$, and if $w\notin\mathfrak{L}(\mathcal{B})$ then $w$ starts with $\heartsuit$ but has exactly $m$ occurrences of $\heartsuit$. Thus, jumping to the first occurrence of a letter in $\Sigma$ incurs a cost of at most $m$, and reading the skipped $\heartsuit$ symbols raises the cost to at most $2m$. From there, $w$ can be read consecutively and be accepted since $w|_{\Sigma}\in\mathfrak{L}(\mathcal{A})$. So again $\mathcal{B}_{\textsc{abs}}(w)\leq 2m$, and $\mathcal{B}$ is $2m$-bounded.
Conversely, if $\mathfrak{L}(\mathcal{A})\neq\Sigma^{*}$, take $x\notin\mathfrak{L}(\mathcal{A})$ such that $x\neq\epsilon$ (see Section A.3 for details regarding this assumption), and consider the word $w=\heartsuit^{m}x$. We then have $w\notin\mathfrak{L}(\mathcal{B})$, and moreover – in order to accept $w$ (if at all possible), $\mathcal{B}$ first needs to jump over the initial $\heartsuit^{m}$, guaranteeing a cost of at least $2m$ ($m$ for the jump and another $m$ to later read the $\heartsuit^{m}$ prefix), and needs at least one more jump to accept $x$, since $x\notin\mathfrak{L}(\mathcal{A})$. Thus, $\mathcal{B}_{\textsc{abs}}(w)>2m$, so $\mathcal{B}$ is not $2m$-bounded. The precise construction and correctness are given in Section A.3. ∎
Lemma 14 shows hardness for fixed $k$, and in particular when $k$ is part of the input. Thus, Univ-Param-Boundedness and Param-Boundedness are also PSPACE-hard, and Univ-$k$-Boundedness is PSPACE-complete.
5 The Reversal Semantics
We now study the reversal semantics. Recall from Definition 4 that for a JFA $\mathcal{A}$ and a word $w$, the cost $\mathcal{A}_{\textsc{rev}}(w)$ is the minimal number of times the jumping head changes “direction” in a jump sequence for which $w$ is accepted.
Consider a word $w$ with $|w|=n$ and a jump sequence $\mathbf{a}=(a_{0},a_{1},a_{2},\ldots,a_{n},a_{n+1})$. We say that an index $1\leq i\leq n$ is a turning index if $a_{i}>a_{i-1}$ and $a_{i}>a_{i+1}$ (i.e., a right-to-left turn) or if $a_{i}<a_{i-1}$ and $a_{i}<a_{i+1}$ (i.e., a left-to-right turn). We denote by $\mathtt{Turn}(\mathbf{a})$ the set of turning indices of $\mathbf{a}$.
For example, consider the jump sequence $(\overset{a_{0}}{0},\overset{a_{1}}{2},\overset{a_{2}}{3},\overset{a_{3}}{5},\overset{a_{4}}{7},\overset{a_{5}}{4},\overset{a_{6}}{1},\overset{a_{7}}{6},\overset{a_{8}}{8})$, then $\mathtt{Turn}(\mathbf{a})=\{4,6\}$. Note that the cost of $w$ is then $\mathcal{A}_{\textsc{rev}}(w)=\min\{|\mathtt{Turn}(\mathbf{a})|\mid w_{\mathbf{a}}\in\mathfrak{L}(\mathcal{A})\}$. Viewed in this manner, we have that $\mathcal{A}_{\textsc{rev}}(w)\leq|w|$, and computing $\mathtt{Turn}(\mathbf{a})$ can be done in polynomial time. Thus, analogously to Theorem 8 we have the following.
Theorem 15 (Reversal Membership is NP-complete).
The problem of deciding, given $\mathcal{A}$ and $k$, whether $\mathcal{A}_{\textsc{rev}}(w)\leq k$ is NP-complete.
Remark 16.
For every jump sequence $\mathbf{a}$ we have that $|\mathtt{Turn}(\mathbf{a})|$ is even, since the head starts at position 00 and ends at $n+1$, where after an odd number of turning points the direction is right-to-left, and hence cannot reach $n+1$.
5.1 Decidability of Boundedness Problems for rev
We begin by characterizing the words accepted using at most $k$ reversals as a shuffle of subwords and reversed-subwords, as follows. Let $x,y\in\Sigma^{*}$ be words, we define their shuffle to be the set of words obtained by interleaving parts of $x$ and parts of $y$. Formally:
[…]
The following lemma states that, intuitively, if $\mathcal{A}_{\textsc{rev}}(w)\leq k$, then $w$ can be decomposed to a shuffle of at most $k+1$ subwords of itself, where all the even ones are reversed (representing the left-reading subwords). See Section A.4 for the proof.
Lemma 17.
Let $k\in\mathbb{N}$. Consider an NFA $\mathcal{A}$ and a word $w\in\Sigma^{*}$. Then $\mathcal{A}_{\textsc{rev}}(w)\leq k$ if and only if there exist words $s_{1},s_{2},\ldots,s_{k+1}\in\Sigma^{*}$ such that the following hold.
  1. 1.
    $s_{1}s_{2}\ldots s_{k+1}\in\mathfrak{L}(\mathcal{A})$.
  2. 2.
[…]
Using the characterization in Lemma 17, we can now construct a corresponding NFA, by intuitively guessing the shuffle decomposition and running copies of $\mathcal{A}$ and its reverse in parallel. See Section A.5 for the proof.
Lemma 18.
Let $k\in\mathbb{N}$ and consider a JFA $\mathcal{A}$. We can effectively construct an NFA $\mathcal{B}$ such that $\mathfrak{L}(\mathcal{B})=\{w\in\Sigma^{*}\mid\mathcal{A}_{\textsc{rev}}(w)\leq k\}$.
The proof of Lemma 18 shows that the size of $\mathcal{B}$ is polynomial in the size of $\mathcal{A}$ and single-exponential in $k$, giving us PSPACE membership for Univ-$k$-Boundedness.
5.2 PSPACE-Hardness of Boundedness for rev
Following a similar scheme to the Absolute Distance Semantics of Section 4, observe that for a word $w\in\Sigma^{*}$ we have that $\mathcal{A}_{\textsc{abs}}(w)=0$ if and only if $w\in\mathfrak{L}(\mathcal{A})$, which implies that Univ-00-Boundedness is PSPACE-hard. Yet again, the challenge is to prove hardness of Univ-$k$-Boundedness for all values of $k$.
Theorem 19.
For rev, Univ-$k$-Boundedness is PSPACE-complete for every $k\in\mathbb{N}$.
Proof.
Membership in PSPACE follows from Lemma 18 and the discussion thereafter. For hardness, we follow the same flow as the proof of Lemma 14, but naturally the reduction itself is different. Specifically, we construct an NFA that must read an expression of the form $(\heartsuit\spadesuit)^{m}$ before its input. This allows us to shuffle the input to the form $\spadesuit^{m}\heartsuit^{m}$, which causes many reversals (see Section A.6). ∎
[…]
6 The Hamming Semantics
Recall from Definition 5 that for a JFA $\mathcal{A}$ and word $w$, the cost $\mathcal{A}_{\textsc{ham}}(w)$ is the minimal Hamming distance between $w$ and $w^{\prime}$ where $w^{\prime}\sim w$ and $w^{\prime}\in\mathfrak{L}(\mathcal{A})$.
Remark 20 (An alternative interpretation of the Hamming Semantics).
We can think of a jumping automaton as accepting a permutation $w^{\prime}$ of the input word $w$. As such, a natural candidate for a quantitative measure is the “distance” of the permutation used to obtain $w^{\prime}$ from the identity (i.e. from $w$). The standard definition for such a distance is the number of transpositions of two indices required to move from one permutation to the other, namely the distance in the Cayley graph for the transpositions generators of $S_{n}$. It is easy to show that in fact, the Hamming distance coincides with this definition.
In Section A.7 we establish the complexity of computing the Hamming measure of a given word.
Theorem 21 (Hamming Membership is NP-complete).
The problem of deciding, given $\mathcal{A}$ and $k\in\mathbb{N}$, whether $\mathcal{A}_{\textsc{ham}}(w)\leq k$ is NP-complete.
Similarly to Sections 4.1 and 5.1, in order to establish the decidability of Univ-Param-Boundedness, we start by constructing an NFA that accepts exactly the words for which $\mathcal{A}_{\textsc{ham}}(w)\leq k$.
Lemma 22.
Let $k\in\mathbb{N}$. We can effectively construct an NFA $\mathcal{B}$ with $\mathfrak{L}(\mathcal{B})=\{w\in\Sigma^{*}\mid\mathcal{A}_{\textsc{ham}}(w)\leq k\}$.
Proof.
Let $k\in\mathbb{N}$. Intuitively, $\mathcal{B}$ works as follows: while reading a word $w$ sequentially, it simulates the run of $\mathcal{A}$, but allows $\mathcal{A}$ to intuitively “swap” the current letter with a (nondeterministically chosen) different one (e.g., the current letter may be $a$ but the run of $\mathcal{A}$ can be simulated on either $a$ or $b$). Then, $\mathcal{B}$ keeps track of the swaps made by counting for each letter $a$ how many times it was swapped by another letter, and how many times another letter was swapped to it. This is done by keeping a counter ranging from $-k$ to $k$, counting the difference between the number of occurrences of each letter in the simulated word versus the actual word. We refer to this value as the balance of the letter. $\mathcal{B}$ also keeps track of the total number of swaps. Then, a run is accepting if at the end of the simulation, the total amount of swaps does not exceed $k$, and if all the letters end up with 00 balance. See Section A.8 for a detailed construction and proof. ∎
An analogous proof to Theorem 12 gives us the following.
Theorem 23.
The following problems are decidable for the ham semantics: $k$-BoundednessParam-BoundednessUniv-$k$-Boundedness, and Univ-Param-Boundedness.
We note that the size of $\mathcal{B}$ constructed in Lemma 22 is polynomial in $k$ and single-exponential in $|\Sigma|$, and therefore when $\Sigma$ is fixed and $k$ is either fixed or given in unary, both Univ-Param-Boundedness and Univ-$k$-Boundedness are in PSPACE.
For a lower bound, we remark that similarly to Section 4.2, it is not hard to prove that Univ-$k$-Boundedness is also PSPACE-hard for every $k$, using relatively similar tricks. However, since Univ-Param-Boundedness is already PSPACE-complete, then Univ-$k$-Boundedness is somewhat redundant. We therefore make do with the trivial lower bound whereby we reduce universality of NFA to Univ-00-Boundedness.
Theorem 24.
For ham, the Univ-Param-Boundedness problem is PSPACE-complete for $k$ encoded in unary and fixed alphabet $\Sigma$.
7 Interplay Between the Semantics
Having established some decidability results, we now turn our attention to the interplay between the different semantics, in the context of boundedness. We show that for a given JFA $\mathcal{A}$, if $\mathcal{A}_{\textsc{abs}}$ is bounded, then so is $\mathcal{A}_{\textsc{ham}}$, and if $\mathcal{A}_{\textsc{ham}}$ is bounded, then so is $\mathcal{A}_{\textsc{rev}}$. We complete the picture by showing that these are the only relationships – we give examples for the remaining cases (see Table 1).
Lemma 25.
Consider a JFA $\mathcal{A}$. If $\mathcal{A}_{\textsc{ham}}$ is bounded, then $\mathcal{A}_{\textsc{rev}}$ is bounded.
Proof.
Consider a word $w\in\Sigma^{*}$, we show that if $\mathcal{A}_{\textsc{ham}}(w)\leq k$ for some $k\in\mathbb{N}$ then $\mathcal{A}_{\textsc{rev}}(w)\leq 3k$. Assume $\mathcal{A}_{\textsc{ham}}(w)\leq k$, then there exists a jump sequence $\mathbf{a}=(a_{0},\ldots,a_{n+1})$ such that $w_{\mathbf{a}}\in\mathfrak{L}(\mathcal{A})$ and $w_{\mathbf{a}}$ differs from $w$ in at most $k$ indices. We claim that we can assume without loss of generality that for every index $i$ such that $w_{a_{i}}=w_{i}$ we have $a_{i}=i$ (i.e., $i$ is a fixed point). Intuitively – there is no point swapping identical letters. Indeed, assume that this is not the case, and further assume that $\mathbf{a}$ has the minimal number of fixed-points among such jump sequences. Thus, there exists some $j$ for which ${a_{j}}\neq j$ but $w_{a_{j}}=w_{j}$. Let $m$ be such that $a_{m}=j$, and consider the jump sequence $\mathbf{a^{\prime}}=(a^{\prime}_{0},\ldots,a^{\prime}_{n+1})$ obtained from $\mathbf{a}$ by composing it with the swap $(a_{j}\ a_{m})$. Then, for every $i\notin\{j,m\}$ we have that $a^{\prime}_{i}=a_{i}$. In addition, $a^{\prime}_{j}=a_{m}=j$ as well as $a^{\prime}_{m}=a_{j}$. In particular, $\mathbf{a^{\prime}}$ has more fixed points than $a$ (exactly those of $\mathbf{a}$ and $j$). However, we claim that $w_{\mathbf{a}}=w_{\mathbf{a^{\prime}}}$. Indeed, the only potentially-problematic coordinates are $a_{j}$ and $a_{m}$. For $j$ we have $w_{a_{j}}=w_{j}=w_{a^{\prime}_{j}}$. and for $m$ we have $w_{a^{\prime}_{m}}=w_{a_{j}}=w_{j}=w_{a_{m}}$. This is a contradiction to $\mathbf{a}$ having a minimal number of fixed points, so we conclude that no such coordinate $a_{j}\neq j$ exists.
Next, observe that $\mathtt{Turn}(\mathbf{a})\subseteq\{i\mid a_{i}\neq i\vee a_{i+1}\neq i+1\vee a_{i-1}\neq i-1\}$. Indeed, if $a_{i-1}=i-1$, $a_{i}=i$ and $a_{i+1}=i+1$ then clearly $i$ is not a turning index. By the property established above, we have that $w_{a_{i}}=w_{i}$, if and only if $a_{i}=i$. It follows that $\mathtt{Turn}(\mathbf{a})\subseteq\{i\mid w_{a_{i}}\neq w_{i}\vee w_{a_{i+1}}\neq w_{i+1}\vee w_{a_{i-1}}\neq w_{i-1}\}$, so $|\mathtt{Turn}(\mathbf{a})|\leq 3k$ (since each index where $w_{\mathbf{a}}\neq w$ is counted at most 3 times [^3] A slightly finer analysis shows that this is in fact at most $2k$, but we are only concerned with boundedness. in the latter set). ∎
Lemma 26.
Consider a JFA $\mathcal{A}$. If $\mathcal{A}_{\textsc{abs}}$ is bounded, then $\mathcal{A}_{\textsc{ham}}$ is bounded.
Proof.
Consider a word $w\in\Sigma^{*}$, we show that if $\mathcal{A}_{\textsc{abs}}(w)\leq k$ for some $k\in\mathbb{N}$ then $\mathcal{A}_{\textsc{ham}}(w)\leq(2k+1)(k+1)$. Assume $\mathcal{A}_{\textsc{abs}}(w)\leq k$, then there exists a jump sequence $\mathbf{a}=(a_{0},\ldots,a_{n+1})$ such that $\llbracket\mathbf{a}\rrbracket\leq k$ and $w_{\mathbf{a}}\in\mathfrak{L}(\mathcal{A})$. In the following we show that $a_{i}=i$ for all but $(2k+1)(k+1)$ indices, i.e., |$|\{i\mid a_{i}\neq i\}|\leq(2k+1)(k+1)$.
It is convenient to think of the jumping head moving according to $\mathbf{a}$ in tandem with a sequential head moving from left to right. Recall that by Lemma 9, for every index $i$ we have that $i-k\leq a_{i}\leq i+k$, i.e. the jumping head stays within distance $k$ from the sequential head.
Consider an index $i$ such that $a_{i}\neq i$ (if there is no such index, we are done). we claim that within at most $2k$ steps, $\mathcal{A}$ performs a jump of cost at least 1 according to $\mathbf{a}$. More precisely, there exists $i+1\leq j\leq i+2k$ such that $|a_{j}-a_{j-1}|>1$. To show this we split to two cases:
  • If $a_{i}>i$, then there exists some $m\leq i$ such that $m$ has not yet been visited according to $\mathbf{a}$ (i.e., by step $i$). Index $m$ must be visited by $a_{i}$ within at most $k$ steps (otherwise it becomes outside the $i-k,i+k$ window around the sequential head), and since $a_{i}>i$, it must perform a “ left jump” of size at least 2 (otherwise it always remains to the right of the sequential reading head).
  • If $a_{i}<i$, then there exists some $m\geq i$ such that $m$ has already been visited by step $i$ according to $\mathbf{a}$. Therefore, within at most $2k$ steps, the jumping head must skip at least over this position (think of $m$ as a hurdle coming toward the jumping head, which must stay within distance $k$ of the sequential head and therefor has to skip over it). Such a jump incurs a cost of at least 1.
Now, let $B=\{i\mid a_{i}\neq i\}$ and assume by way of contradiction that $|B|>(2k+1)(k+1)$. By the above, for every $i\in B$, within $2k$ steps the run incurs a cost of at least 1. While some of these intervals of $2k$ steps may overlap, we can still find at least $k+1$ such disjoint segments (indeed, every $i\in B$ can cause an overlap with at most $2k$ other indices). More precisely, there are $i_{1}<i_{2}<\ldots<i_{k+1}$ in $B$ such that $i_{j}>i_{j-1}+2k$ for all $j$, and therefore each of the costs incurred within $2k$ steps of visiting $i_{j}$ is independent of the others. This, however, implies that $\llbracket\mathbf{a}\rrbracket\geq k+1$, which is a contradiction, so $|B|\leq(2k+1)(k+1)$.
It now follows that $\mathcal{A}_{\textsc{ham}}(w)=|\{i\mid w_{a_{i}}\neq w_{i}\}|\leq|\{i\mid a_{i}\neq i\}|\leq(2k+1)(k+1)$
Combining Lemmas 25 and 26, we have the following.
Corollary 27.
If abs is bounded, then so is rev.
We proceed to show that no other implication holds with regard to boundedness, by demonstrating languages for each possible choice of bounded/unbounded semantics (c.f. Remark 6). The examples are summarized in Table 1, and are proved below.
abs ham rev Language
Bounded Bounded Bounded $(a+b)^{*}$
Unbounded Bounded Bounded $(a+b)^{*}a$
Unbounded Unbounded Bounded $a^{*}b^{*}$
Unbounded Unbounded Unbounded $(ab)^{*}$
Table 1: Examples for every possible combination of bounded/unbounded semantics. The languages are given by regular expressions (e.g., $(a+b)^{*}a$ is the languages of words that end with $a$.)
Example 28.
The language $(a+b)^{*}$ is bounded in all semantics. This is trivial, since every word is accepted, and in particular has cost 00 in all semantics.
Example 29.
The language $(a+b)^{*}a$ is bounded in the ham and rev semantics, but unbounded in abs. Indeed, let $\mathcal{A}$ be an NFA such that $\mathfrak{L}(\mathcal{A})=(a+b)^{*}a$ and consider a word $w\in\mathfrak{J}(\mathcal{A})$, then $w$ has at least one occurrence of $a$ at some index $i$. Then, for the jumping sequence $\mathbf{a}=(0,1,2,\ldots,i-1,n,i+1,\ldots n-1,i,n+1)$ we have that $w_{\mathbf{a}}\in\mathfrak{L}(\mathcal{A})$. Observe that $d_{H}(w_{\mathbf{a}},w)\leq 2$ (since $w_{\mathbf{a}}$ differs from $w$ only in indices $i$ and $n$), and $\mathtt{Turn}(\mathbf{a})\subseteq\{i,n\}$, so $\mathcal{A}_{\textsc{ham}}(w)\leq 2$ and $\mathcal{A}_{\textsc{rev}}(w)\leq 2$.
For abs, however, consider the word $ab^{n}$ for every $n\in\mathbb{N}$. Since the letter $a$ must be read last, then in any jumping sequence accepting the word, there is a point where the jumping head is at index $n$ and the sequential head is at position $1$. By Lemma 9, it follows that $\mathcal{A}_{\textsc{abs}}(w)\geq n-1$, and by increasing $n$, we have that $\mathcal{A}_{\textsc{abs}}$ is unbounded.
Example 30.
The language $a^{*}b^{*}$ is bounded in the rev semantics, but unbounded in ham and abs. Indeed, let $\mathcal{A}$ be an NFA such that $\mathfrak{L}(\mathcal{A})=a^{*}b^{*}$ and consider a word $w\in\mathfrak{J}(\mathcal{A})$, and denote by $i_{1}<i_{2}\ldots<i_{k}$ the indices of $a$’s in $w$ in increasing order, and by $j_{1}>j_{2}>\ldots>j_{n-k}$ the indices of $b$’s in decreasing order. Then, for the jumping sequence $\mathbf{a}=(i_{1},\ldots,i_{k},j_{1},\ldots j_{n-k},n+1)$ we have that $w_{\mathbf{a}}\in\mathfrak{L}(\mathcal{A})$, and $\mathcal{A}_{\textsc{rev}}(w)\leq 2$ (since the jumping head goes right reading all the $a$’s, then left reading all the $b$’s, then jumps to $n+1$).
For ham, consider the word $w=b^{n}a^{n}$ for every $n\in\mathbb{N}$. The only permutation of $w$ that is accepted in $\mathfrak{L}(\mathcal{A})$ is $w^{\prime}=a^{n}b^{n}$, and $d_{H}(w,w^{\prime})=n$, so $\mathcal{A}_{\textsc{ham}}$ is unbounded. By Lemma 26 it follows that $\mathcal{A}_{\textsc{abs}}$ is also unbounded.
Example 31.
The language $(ab)^{*}$ is unbounded in all the semantics. Indeed, let $\mathcal{A}$ be an NFA such that $\mathfrak{L}(\mathcal{A})=(ab)^{*}$, then by Lemmas 25 and 27 it suffices to show that $\mathcal{A}_{\textsc{rev}}$ is unbounded.
Consider the word $w=b^{n}a^{n}$ for every $n\in\mathbb{N}$, and let $\mathbf{a}=(a_{0},a_{1},\ldots,a_{2n},a_{2n+1})$ such that $w_{\mathbf{a}}\in(ab)^{*}$, then for every odd $i\leq 2n$ we have $a_{i}\in\{n+1,\ldots,2n\}$ and for every even $i\leq 2n$ we have $a_{i}\in\{1,\ldots,n\}$. In particular, every index $1\leq i\leq 2n$ is a turning point, so $\mathcal{A}_{\textsc{rev}}(w)=2n$, and $\mathcal{A}_{\textsc{rev}}$ is unbounded.
8 Discussion and Future Work
Quantitative semantics are often defined by externally adding some quantities (e.g., weights) to a finite-state model, usually with the intention of explicitly reasoning about some unbounded domain. It is rare and pleasing when quantitative semantics arise naturally from a Boolean model. In this work, we study three such semantics. Curiously, despite the semantics being intuitively unrelated, it turns out that they give rise to interesting interplay (see Section 7).
We argue that Boundedness is a fundamental decision problem for the semantics we introduce, as it measures whether one can make do with a certain budget for jumping. An open question left in this research is existentially-quantified boundedness: whether there exists some bound $k$ for which $\mathcal{A}_{\textsc{sem}}$ is $k$-bounded. This problem seems technically challenging, as in order to establish its decidability, we would need to upper-bound the minimal $k$ for which the automaton is $k$-bounded, if it exists. The difficulty arises from two fronts: first, standard methods for showing such bounds involve some pumping argument. However, the presence of permutations makes existing techniques inapplicable. We expect that a new toolbox is needed to give such arguments. Second, the constructions we present for Univ-Param-Boundedness in the various semantics seem like the natural approach to take. Therefore, a sensible direction for the existential case is to analyze these constructions with a parametric $k$. The systems obtained this way, however, do not fall into (generally) decidable classes. For example, in the ham semantics, using a parametric $k$ we can construct a labelled VASS. But the latter do not admit decidable properties for the corresponding boundedness problem.
We remark on one fragment that can be shown to be decidable: consider a setting where the jumps are restricted to swapping disjoint pairs of adjacent letters, each incurring a cost of 1. Then, the JFA can be translated to a weighted automaton, whose boundedness problem is decidable by [16, 19]. We remark that the latter decidability is a very involved result. This suggests (but by no means proves) that boundedness may be a difficult problem.
Appendix A Proofs
A.1 Proof of Lemma 9
[…]
$\sum_{i=j+1}^{n+1}\llbracket a_{i}-a_{i-1}\rrbracket=\left(\sum_{i=j+1}^{n+1}|a_{i}-a_{i-1}|\right)-(n+1-(j+1)+1)=\left(\sum_{i=j+1}^{n+1}|a_{i}-a_{i-1}|\right)-(n+1-j)$
Similarly to the previous case, from the triangle inequality we have
$\left(\sum_{i=j+1}^{n+1}|a_{i}-a_{i-1}|\right)-(n+1-j)\geq|a_{n+1}-a_{j}|-(n+1-j)=n+1-a_{j}-(n+1-j)=j-a_{j}=|a_{j}-j|$
where we use the fact that $a_{n+1}=n+1>a_{j}$, and our assumption that $a_{j}<j$. This again concludes that $\llbracket\mathbf{a}\rrbracket\geq|a_{j}-j|$. ∎
A.2 Proof of Theorem 12
Consider a JFA $\mathcal{A}$ and $k\in\mathbb{N}$ ($k$ is either fixed or given as input, which does not effect decidability), and let $\mathcal{B}$ be the NFA constructed as per Lemma 11. In order to decide Univ-$k$-Boundedness and Univ-Param-Boundedness, observe that $\mathcal{A}_{\textsc{abs}}(w)\leq k$ for every word $w\in\Sigma^{*}$ if and only if $\mathfrak{L}(\mathcal{B})=\Sigma^{*}$. Since the latter is decidable for NFA, we have decidability.
Similarly, in order to decide $k$-Boundedness and Univ-Param-Boundedness, observe that $\mathcal{A}_{\textsc{abs}}(w)\leq k$ for every word $w\in\mathfrak{J}(\mathcal{A})$ if and only if $\mathfrak{J}(\mathcal{A})\subseteq\mathfrak{L}(\mathcal{B})$. We can decide whether the latter holds by constructing the PA (¯$(\overline{\mathcal{B}},\mathcal{A})$ where $\overline{\mathcal{B}}$ is an NFA for the complement of $\mathfrak{L}(\mathcal{B})$, and checking emptiness. Since emptiness for PA is decidable [17], we conclude decidability. ∎
A.3 Proof of Lemma 14
In order for the reduction to work both for Univ-$k$-Boundedness and $k$-Boundedness, we start with an initial transformation of the given NFA $\mathcal{A}$ to an NFA $\mathcal{A}^{\prime}$ as follows. Given $\mathcal{A}$, we introduce a fresh symbol \$currency-dollar\\$\$ to its alphabet, and modify it so that reading \$currency-dollar\\$\$ from every state can either stay at the same state, or transition to a new accepting state $q_{\$}$. From $q_{\$}$ no letters can be read. $\mathcal{A}^{\prime}$ then satisfies the following property: if $L(\mathcal{A})=\Sigma^{*}$ then $L(\mathcal{A}^{\prime})=(\Sigma\cup\{\$\})^{*}$, and if $L(\mathcal{A})\neq\Sigma^{*}$ then there exists a word $x\notin L(\mathcal{A}^{\prime})$ such that $x\in\mathfrak{J}(\mathcal{A}^{\prime})$. Indeed, for every (non-empty) word $x\notin L(\mathcal{A})$ we have that \$$\$x\notin L(\mathcal{A}^{\prime})$ but $x\$\in L(\mathcal{A}^{\prime})$. We henceforth identify $\mathcal{A}$ with $\mathcal{A}^{\prime}$, and use this property in the proof.
We now turn to the details of the reduction, differing from the intuition in the main body mainly by the pesky presence of $\epsilon$. Given $\mathcal{A}$, we first check whether $\epsilon\in\mathfrak{L}(\mathcal{A})$ (i.e., we check whether $Q_{0}\cap\alpha\neq\emptyset$). If $\epsilon\notin\mathfrak{L}(\mathcal{A})$, then $\mathfrak{L}(\mathcal{A})\neq\Sigma^{*}$ and we output some fixed unbounded automaton $\mathcal{B}$ (e.g., as in Examples 1 and 31). Otherwise, we proceed to construct $\mathcal{B}$ such that $w\in\mathfrak{L}(\mathcal{B})$ if and only if either Conditions 1 or 2 above hold, or:
  1. 3.
    $w=\heartsuit^{m}$ (think of this as an exception to Condition 2 where $w|_{\Sigma}=\epsilon$).
Constructing $\mathcal{B}$ from $\mathcal{A}$ is straightforward by taking roughly $m$ copies of $\mathcal{A}$ to track the number of $\heartsuit$ in the word. In particular, the reduction is in polynomial time.
We claim that $\mathfrak{L}(\mathcal{A})=\Sigma^{*}$ if and only if $\mathcal{B}$ is $2m$-bounded. For the first direction, assume $\mathfrak{L}(\mathcal{A})=\Sigma^{*}$. As mentioned above, for every $w\in(\Sigma\cup\{\heartsuit\})^{*}$ if $w\in\mathfrak{L}(\mathcal{B})$ then $\mathcal{B}_{\textsc{abs}}(w)=0\leq 2m$. If $w\notin\mathfrak{L}(\mathcal{B})$ then $w$ starts with $\heartsuit$ but has exactly $m$ occurrences of $\heartsuit$, and in addition, $w\neq\heartsuit^{m}$ so the $m+1$-th letter of $w$ is in $\Sigma$. Then, the following jump sequence $\mathbf{a}$ causes $\mathcal{B}$ to accept $w_{\mathbf{a}}$:
$\mathbf{a}=(0,m+1,1,2,\ldots,m,m+2,m+3,\ldots|w|,|w|+1)$
Indeed, $w_{\mathbf{a}}$ does not start with $\heartsuit$, has exactly $m$ occurrences of $\heartsuit$, and $w|_{\Sigma}\in\mathfrak{L}(\mathcal{A})=\Sigma^{*}$, so Condition 2 holds. Finally, note that $\llbracket\mathbf{a}\rrbracket=2m$ (since the only non-zero jumps are 00 to $m+1$, $m+1$ to $1$, and $m$ to $m+2$).
For the converse, assume $\mathfrak{L}(\mathcal{A})\neq\Sigma^{*}$. If $\epsilon\notin\mathfrak{L}(\mathcal{A})$, then correctness follows by the treatment of this case above. Otherwise, let $x\notin\mathfrak{L}(\mathcal{A})$ with $x\neq\epsilon$. We can assume by the modification above that $x\in\mathfrak{J}(\mathcal{A})$. and consider $w=\heartsuit^{m}x$. We claim that $\mathcal{B}_{\textsc{abs}}(w)>2m$. Indeed, $w\notin\mathfrak{L}(\mathcal{B})$, and therefore $\mathcal{B}_{\textsc{abs}}(w)>0$. If $\mathcal{B}_{\textsc{abs}}(w)=\infty$ then we are done. Otherwise, let $\mathbf{a}$ be a jump sequence such that $w_{\mathbf{a}}\in\mathfrak{L}(\mathcal{B})$. Then $a_{1}\geq m+1$, contributing a cost of at least $m$ to $\mathbf{a}$. It can be easily seen by induction over $m$ that in order for $\mathbf{a}$ to cover the entries $1,\ldots,m$ starting at position $m+1$ and ending at position $m+2$, it requires cost of at least another $m$. Then, however, in order for $w_{\mathbf{a}}$ to be accepted by $\mathcal{B}$, it must hold that $w_{\mathbf{a}}|_{\Sigma}\neq x$, so $\mathbf{a}$ is not the identity starting from $m+1$. It therefore has an additional cost of at least $1$. Thus, $\llbracket\mathbf{a}\rrbracket>2m$. In particular, $\mathcal{B}_{\textsc{abs}}(w)>2m$, so $\mathcal{B}$ is not $2m$-bounded. Note that $\mathcal{B}_{\textsc{abs}}(w)<\infty$, since $x\in\mathfrak{J}(\mathcal{A})$. Thus, we are done both in the universal and non-universal setting. ∎
A.4 Proof of Lemma 17
For the first direction, assume $\mathcal{A}_{\textsc{rev}}(w)\leq k$, so there exists a jump vector $\mathbf{a}$ such that $|\mathtt{Turn}(\mathbf{a})|\leq k$ and $w_{\mathbf{a}}\in\mathfrak{L}(\mathcal{A})$. Write $\mathtt{Turn}(\mathbf{a})=\{i_{1},i_{2},\ldots,i_{\ell}\}$ where $i_{1}<i_{2}<\ldots<i_{\ell}$, and set $i_{0}=0$ and $i_{\ell+1}=n+1$. Then, for every $11\leq j\leq\ell+$ 1, the $j$-th turning subword is $s_{j}=w_{a_{i_{j-1}}}w_{a_{i_{j-1}+1}}\cdots w_{a_{i_{j}}}$ (for $j=\ell+1$ we end with $w_{a_{i_{\ell+1}-1}}$). If $\ell+1<k+1$ we define the remaining subwords $s_{\ell+2},\ldots,s_{k+1}$ to be $\epsilon$. To avoid cumbersome indexing, we assume $\ell+1=k+1$ in the following.
[…]
A.5 Proof of Lemma 18
The overall plan is to construct $\mathcal{B}$ so that it captures the conditions in Lemma 17. Intuitively, $\mathcal{B}$ keeps track of $k+1$ coordinates, each corresponding to a turning subword (that are nondeterministically constructed). The odd coordinates simulate the behavior of $\mathcal{A}$, whereas the even ones simulate the reverse of $\mathcal{A}$. In addition $\mathcal{B}$ checks (using its initial and accepting states) that the runs on the subwords can be correctly concatenated. We proceed with the precise details.
Denote $\mathcal{A}=\langle\Sigma,Q,\delta,Q_{0},F\rangle$. We construct $\mathcal{B}=\langle\Sigma,Q^{\prime},\delta^{\prime},Q^{\prime}_{0},F^{\prime}\rangle$ as follows. $Q^{\prime}=Q^{k+1}$, and the initial and final states are:
$Q^{\prime}_{0}=\{(q_{1},q_{2},\ldots q_{k+1})\ |\ q_{1}\in Q_{0}\wedge q_{i}=q_{i+1}\text{ for all even }i\}$
$F^{\prime}=\{(q_{1},q_{2},\ldots q_{k+1})\ |\ q_{k+1}\in F\wedge q_{i}=q_{i+1}\text{ for all odd }i\}$
For the transition function, we have that $(q^{\prime}_{0},q^{\prime}_{1},\ldots q^{\prime}_{k})\in\delta^{\prime}((q_{0},q_{1},\ldots q_{k}),\sigma)$ if and only if there exists a single $11\leq j\leq k+$ 1 such that $q^{\prime}_{j}\in\delta(q_{j},\sigma)$ if $j$ is odd, and $q_{j}\in\delta(q^{\prime}_{j},\sigma)$ if $j$ is even. In addition, for every $i\neq j$, it holds that $q^{\prime}_{i}=q_{i}$.
We turn to show the correctness of $\mathcal{B}$. Consider an accepting run $\rho$ of $\mathcal{B}$ on some word. Then $\rho$ starts at state $(q_{1},q_{2},\ldots,q_{k+1})\in Q^{\prime}_{0}$ and ends at state $(s_{1},s_{2},\ldots,s_{k+1})\in F^{\prime}$. By the definition of $\delta^{\prime}$, we can split $\rho$ according to which component “progresses” in each transition, so that $\rho$ can be written as a shuffle of run $\rho^{1},\ldots,\rho^{k+1}$ where $\rho^{i}$ leads from $q_{i}$ to $s_{i}$ in $\mathcal{A}$ if $i$ is odd, and $\rho^{i}$ leads from $q_{i}$ to $s_{i}$ in the reverse of $\mathcal{A}$ if $i$ is even. The latter is equivalent to $(\rho^{i})^{R}$ (i.e., the reverse run of $\rho^{i}$) leading from $s_{i}$ to $q_{i}$ in $\mathcal{A}$ if $i$ is even.
We now observe that these runs can be concatenated as follows: Recall that $q_{1}\in Q_{0}$ (by the definition of $Q^{\prime}_{0}$). Then, $\rho^{1}$ leads from $q_{1}$ to $s_{1}$ in $\mathcal{A}$. By the definition of $F$ we have $s_{1}=s_{2}$, and $(\rho^{2})^{R}$ leads from $s_{2}$ to $q_{2}$ in $\mathcal{A}$. Therefore, $\rho^{1}(\rho^{2})^{R}$ leads from $q_{1}$ to $q_{2}$ in $\mathcal{A}$. Continuing in the same fashion, we have $q_{2}=q_{3}$, and $\rho^{3}$ leading from $q_{3}$ to $s_{3}$, and so on up to $s_{k+1}$.
Thus, we have that $\rho^{1}(\rho^{2})^{R}\rho_{3}\cdots(\rho^{k})^{R}\rho^{k+1}$ is an accepting run of $\mathcal{A}$.
By identifying each accepting run $\rho^{i}$ with the subword it induces, we have that $w\in\mathfrak{L}(\mathcal{B})$ if and only if there are words $s_{1},\ldots,s_{k+1}$ such that the two conditions in Lemma 17 are satisfied. ∎
A.6 Proof of Theorem 19
By Remark 16 we can assume without loss of generality that $k$ is even, and we denote $k=2m$. We reduce the universality problem for NFAs to the Univ-$2m$-Boundedness problem. Consider an NFA $\mathcal{A}=\langle Q,\Sigma,\delta,Q_{0},\alpha\rangle$, and let $\heartsuit,\spadesuit\notin\Sigma$ be fresh symbols. We first check whether $\epsilon\in\mathfrak{L}(\mathcal{A})$. If $\epsilon\notin\mathfrak{L}(\mathcal{A})$, then $\mathfrak{L}(\mathcal{A})\neq\Sigma^{*}$ and we output some fixed unbounded automaton $\mathcal{B}$ (e.g., as in Examples 1 and 31).
Otherwise, we obtain from $\mathcal{A}$ an NFA $\mathcal{B}$ over the alphabet $\Sigma\cup\{\heartsuit,\spadesuit\}$ such that $w\in\mathfrak{L}(\mathcal{B})$ if and only if the following hold:
  1. 1.
    Either $w$ does not contain exactly $m$ occurrences of $\heartsuit$ and of $\spadesuit$, or
  2. 2.
    $w=(\heartsuit\spadesuit)^{m}x$ where $x\in\mathfrak{L}(\mathcal{A})$ (in particular $x\in\Sigma^{*}$).
Constructing $\mathcal{B}$ from $\mathcal{A}$ is straightforward as the union of two components: one that accepts words that satisfy Condition 1 (using $12m+$ 1 states) and one for Condition $2$, which prepends to $\mathcal{A}$ a component with $2m$ states accepting $(\heartsuit\spadesuit)^{m}$. In particular, the reduction is in polynomial time.
We then have the following: if $\mathfrak{L}(\mathcal{A})=\Sigma^{*}$, then for every $w\in(\Sigma\cup\{\heartsuit,\heartsuit\})^{*}$, if $w$ satisfies Condition 1, then $\mathcal{B}_{\textsc{rev}}(w)=0$. Otherwise, $w$ has exactly $m$ occurrences of $\heartsuit$ and of $\spadesuit$. Denote the indices of $\heartsuit$ by $i_{1}<i_{2}<\ldots<i_{m}$ and of $\spadesuit$ by $j_{1}<j_{2}<\ldots<j_{m}$. Also denote by $t_{0}<t_{1}<\ldots<t_{r}$ the remaining indices of $w$. Then, consider the jump sequence
$\mathbf{a}=(0,i_{1},j_{1},i_{2},j_{2},\ldots,i_{m},j_{m},t_{0},t_{1},\ldots,t_{r},n+1)$
We claim that $w_{\mathbf{a}}\in\mathfrak{L}(\mathcal{B})$ by Condition 2. Indeed, $w$ starts with $(\heartsuit\spadesuit)^{m}$, followed by letters in $\Sigma$ composing a word $x$. Since $x\in\mathfrak{L}(\mathcal{A})=\Sigma^{*}$, we have that Condition 2 holds. In addition, observe that since $t_{0}<t_{1}<\ldots<t_{r}<n+1$, then $\mathtt{Turn}(\mathbf{a})\subseteq\{i_{1},j_{1},\ldots,i_{m},j_{m},t_{0}\}$, and in particular $|\mathtt{Turn}(\mathbf{a})|\leq 2m+1$. Moreover, by Remark 16 we know that $|\mathtt{Turn}(\mathbf{a})|$ is even, so in fact $|\mathtt{Turn}(\mathbf{a})|\leq 2m=k$. We conclude that $\mathcal{B}_{\textsc{rev}}(w)\leq k$, so $\mathcal{B}$ is $k$-bounded.
Conversely, if $\mathfrak{L}(\mathcal{A})\neq\Sigma^{*}$, take $x\notin\mathfrak{L}(\mathcal{A})$ such that $x\neq\epsilon$ (which exists since we checked above that $\epsilon\in\mathfrak{L}(\mathcal{A})$). Consider the word $w=\spadesuit^{m}\heartsuit^{m}x$, then have $w\notin\mathfrak{L}(\mathcal{B})$. We claim that $\mathcal{B}_{\textsc{rev}}(w)>2m$. Indeed, if there exists $\mathbf{a}$ such that $w_{\mathbf{a}}\in\mathfrak{L}(\mathcal{B})$, then since $w$ has exactly $m$ occurrences of $\spadesuit$ and of $\heartsuit$, it must be accepted by Condition 2. By the structure of $w$, the jump sequence $\mathbf{a}$ needs to permute $\spadesuit^{m}\heartsuit^{m}$ into $(\heartsuit\spadesuit)^{m}$. Intuitively, this means that the head must jump “back and forth” for $2m$ steps. More precisely, for every $i\in\{1,\ldots,|w|\}$ it holds that
$a_{i}\in\begin{cases}\{m+1,\ldots,2m\}&i\leq 2m\text{ is odd}\\ \{1,\ldots,m\}&i\leq 2m\text{ is even}\\ \{2m+1,\ldots,|w|\}&i>2m\\ \end{cases}$
In particular, $\{1,\ldots,2m\}\subseteq\mathtt{Turn}(\mathbf{a})$. Observe that the remaining suffix of $w_{\mathbf{a}}$ starting at $12m+$ 1 cannot be $x$, since $x\notin\mathfrak{L}{\mathcal{A}}$, so $\mathbf{a}$ is not the identity starting from $12m+$ 1. It therefore has an additional reversal cost of at least $1$. Thus, $|\mathtt{Turn}(\mathbf{a})|>2m$. In particular, $\mathcal{B}_{\textsc{rev}}(w)>2m$, so $\mathcal{B}$ is not $2m$-bounded, and we are done. ∎
A.7 Proof of Theorem 21
By definition we have that $\mathcal{A}_{\textsc{ham}}(w)\leq|w|$ for every word $w$. Thus, in order to decide whether $\mathcal{A}_{\textsc{ham}}(w)\leq k$ we can nondeterministically guess a permutation $w^{\prime}\sim w$ and verify that $w^{\prime}\in\mathfrak{L}(\mathcal{A})$ and that $d_{H}(w,w^{\prime})\leq k$. Both conditions are computable in polynomial time. Therefore, the problem is in NP.
Hardness follows (similarly to the proof of Theorem 8) by reduction from membership in JFA, noting that $w\in\mathfrak{J}(\mathcal{A})$ if and only if $\mathcal{A}_{\textsc{ham}}(w)\leq|w|$. ∎
A.8 Proof ofLemma 22
We now turn to the formal details. Recall that $\mathcal{A}=\langle\Sigma,Q,\delta,Q_{0},\alpha\rangle$. We define $\mathcal{B}=\langle\Sigma,Q^{\prime},\delta^{\prime},Q_{0}^{\prime},\beta\rangle$.
The state space of $\mathcal{B}$ is $Q^{\prime}=Q\times\{-k,\ldots,k\}^{\Sigma}\times\{0,\ldots,k\}$. We denote a state of $\mathcal{B}$ by $(q,f,c)$ where $q\in Q$ is the current state of $\mathcal{A}$, $f:\Sigma\to\{-k,\ldots,k\}$ describes for each letter its balance and $c\in\{0,\ldots,k\}$ is the total number of swaps thus far.
The initial states of $\mathcal{B}$ are $Q^{\prime}_{0}=\{(q,f,0)\mid q\in Q_{0}\wedge f(\sigma)=0\text{ for all }\sigma\in\Sigma\}$. That is, we start in an initial state of $\mathcal{A}$ with balance and total cost of 00. The transition function is defined as follows. Consider a state $(q,f,c)$ and a letter $\sigma\in\Sigma$, then $(q^{\prime},f^{\prime},c^{\prime})\in\delta^{\prime}((q,f,c),\sigma)$ if and only if either $q^{\prime}=\delta(q,\sigma)$ and $f^{\prime}=f$ and $c^{\prime}=c$, or there exists $\tau\in\Sigma$, $\tau\neq\sigma$ such that $q^{\prime}\in\delta(q,\tau)$, $c^{\prime}=c+1$, $f^{\prime}(\sigma)=f(\sigma)-1$, and $f^{\prime}(\tau)=f(\tau)+1$. That is, in each transition we either read the current letter $\sigma$, or swap for a letter $\tau$ and update the balances accordingly.
Finally, the accepting states of $\mathcal{B}$ are $\beta=\{(q,f,c)\mid q\in\alpha\wedge f(\sigma)=0\text{ for all }\sigma\in\Sigma\}$.
In order to establish correctness, we observe that every run of $\mathcal{B}$ on a word $w$ induces a word $w^{\prime}$ (with the nondeterministically guessed letters) such that along the run the components $f$ and $c$ of the states track the swaps made between $w$ and $w^{\prime}$. In particular, $c$ keeps track of the number of total swaps, and $\sum_{\sigma\in\Sigma}f(\sigma)=0$. Moreover, for every word $\sigma$, the value $f(\sigma)$ is exactly the number of times $\sigma$ was read in $w^{\prime}$ minus the number of times $\sigma$ was read in $w$.
Since $\mathcal{B}$ accepts a word only if $f\equiv 0$ at the last state, it follows that $\mathcal{B}$ accepts if and only if $w^{\prime}\sim w$, and the run of $\mathcal{A}$ on $w^{\prime}$ is accepting. Finally, since $c$ is bounded by $k$ and is increased upon every swap, then $w$ is accepted if and only if its cost is at most $k$. Note that since $c$ is increased upon each swap, then limiting the image of $f$ to values in $\{-k,\ldots,k\}$ does not pose a restriction, as they cannot go beyond these bounds without $c$ going beyond the bound $k$ as well. ∎